OpenStudy (nbarrera):

Which number is between 0.5 and 5/8? A.8/28 B.1/4 C.19/28 D.15/28<= my answer

OpenStudy (anonymous):

C

OpenStudy (anonymous):

0.5 is.. 0.50 So what ever number is between 0.50 and 0.62... 5/8 is..= 5/8 = 5 ÷ 8 = 0.625 A) 8/28 = 2/7 = 2 ÷ 7 = 0.285714 B) 1/4 = 1/4 = 1 ÷ 4 = 0.25 c) 19/28 = 19/28 = 19 ÷ 28 = 0.678571 D) 15/28 = 15/28 = 15 ÷ 28 = 0.535714 So yes your right about it being D.

OpenStudy (nbarrera):

Ok thanks

OpenStudy (anonymous):

If you got anymore. Just tag me in them. And i'll try my hardest to help you out.

OpenStudy (nbarrera):

no that all

OpenStudy (anonymous):

okay(: Good luck.

OpenStudy (nbarrera):

Yay got all the questions right